Suppose the call money rate is 4.5 percent, and you pay a spread of 2.5 percent over that. You buy 900 shares of stock at $46 per share. You put up $24,840. One year later, the stock is selling for $60 per share and you close out your position. What is your return assuming a dividend of $0.23 per share is paid? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.) Rate of return
Answer #1
Initial investment $ 24,840.00
Amount borrowed $ 16,560.00 =900*46-24840
Interest expense $    1,159.20 =16560*(4.5%+2.5%)
Dollar return $ 11,647.80 =900*(60+0.23-46)-1159.2
Rate of return 46.89% =11647.8/24840